Equal volumes, 50.0 ml of 0.54897 M solution of HCl and 0.9573 M NaOH solutions having an initial temperature of 18.5 degree Celsius react in a coffee cup calorimeter. The resultant solution records a temperature of 22.9 degree Celsius. since the solutions are mostly water, the solutions are assumed to have a density of 1.0g/ml and a specific heat of 4.184 j/g degree Celsius. The heat capacity of the calorimeter is 1.0*10^1 j/degree Celsius. Calculate hrxn(kJ/mol)

1
Expert's answer
2021-11-12T09:36:02-0500

Number of moles of HCl and NaOH added= 1000


MV= 1/1000×50

=0.05

Mass of mixture =V×d=100×=100g

Heat evolved, q=msΔT=100×4.18×(27.5−21.0)

=100×4.18×6.5J=2717J=2.717 kJ


(b) The involved reaction is:

HCl(aq.)+NaOH(aq.)→NaCl(aq.)+H2O

ΔH= Heat evolved per mol

= 2.717÷0.05

=54.34 kJ